OpenStudy (anonymous):

Using the first 4 terms of the Maclaurin series, what is the value of e^1.4?

OpenStudy (anonymous):

maclaurin series just taylor yes?

OpenStudy (anonymous):

\[e^x=1+x+\frac{x^2}{2}+\frac{x^3}{3!}+...\]

OpenStudy (anonymous):

replace x by 1.4 get 3.837333333